OpenStudy (anonymous):

two right circular cones have equal bases. the altitude of one cone is four times the diameter of its base while the other cone has an altitude equal to is diameter. Find the dimensions of the larger cone if the difference of their volumes is 16pi cubic units.

OpenStudy (anonymous):

@sillynikki1

OpenStudy (anonymous):

Suppose h1=4D h2=D V1-V2=16pi 1/3.pi.r^2.h1-1/3.pi.r^2.h2=16pi 1/3.pi.r^2(h1-h2)=16pi r^2(4D-D)=48 3D.r^2=48 D=2r, so: 6r^3=48 r^3=8 r=2 So, the dimension of the larger cone is: h1=4D=4.2r=4.2.2=16 and r=2 The dimension is a cone with height 16 unit and radius 2 unit

OpenStudy (anonymous):

Not understandable?